### v3

# Pre-Chirp-Domain Index Modulation for Full-Diversity Affine Frequency Division Multiplexing towards 6G

Abstract
As a superior multicarrier technique utilizing chirp signals for high-mobility communications, affine frequency division multiplexing (AFDM) is envisioned to be a promising candidate for sixth-generation (6G) wireless networks. AFDM is based on the discrete affine Fourier transform (DAFT) with two adjustable parameters of the chirp signals, termed the pre-chirp and post-chirp parameters, respectively. Whilst the post-chirp parameter complies with stringent constraints to combat the time-frequency doubly selective channel fading, we show that the pre-chirp counterpart can be flexibly manipulated for an additional degree of freedom. Therefore, this paper proposes a novel AFDM scheme with the pre-chirp index modulation (PIM) philosophy (AFDM-PIM), which can implicitly convey extra information bits through dynamic pre-chirp parameter assignment, thus enhancing both spectral and energy efficiency. Specifically, we first demonstrate that the subcarrier orthogonality is still maintained by applying distinct pre-chirp parameters to various subcarriers in the AFDM modulation process. Inspired by this property, we allow each AFDM subcarrier to carry a unique pre-chirp signal according to the incoming bits. By such an arrangement, extra bits can be embedded into the index patterns of pre-chirp parameter assignment without additional energy consumption. We derive asymptotically tight upper bounds on the average bit error probability (BEP) of the proposed schemes with the maximum-likelihood detection, and validate that the proposed AFDM-PIM can achieve full diversity under doubly dispersive channels. Based on the derived result, we further propose an optimal pre-chirp alphabet design to enhance the bit error rate (BER) performance via intelligent optimization algorithms. Simulation results demonstrate that the proposed AFDM-PIM outperforms the classical benchmarks.
Index Terms:
Index modulation (IM), affine frequency division multiplexing (AFDM), discrete affine Fourier transform (DAFT), doubly dispersive channel.
I Introduction
The beyond fifth-generation (B5G) and sixth-generation (6G) wireless networks are envisioned to deliver ultra-reliable, high data rate, and low-latency communications for high–speed mobile scenarios, including low-earth-orbit (LEO) satellite, high-mobility railway, unmanned aerial vehicles (UAV) and vehicle-to-vehicle (V2V) communications [2, 3, 4, 5]. These scenarios inevitably suffer from severe Doppler shifts, which can cause time-frequency doubly selective fading (i.e., doubly dispersive channel) by involving the multi-path effects. This makes the existing modulation formats, like the main-stream orthogonal frequency division multiplexing (OFDM) in 4G/5G standards, no longer suitable for next-generation networks [6]. Consequently, it is crucial to develop new waveforms for next-generation communication networks to adapt to the doubly selective channel.
To date, several novel modulation techniques have been designed to combat the time-frequency doubly selective fading, such as orthogonal time-frequency space (OTFS) [7, 8, 9] and orthogonal chirp-division multiplexing (OCDM) [10, 11, 12]. OTFS modulates information in the delay-Doppler (DD) domain using the inverse symplectic finite Fourier transform (ISFFT), which enables the transmission symbols to be multiplexed across the entire time-frequency domain [13, 14, 15]. OCDM utilizes a series of orthogonal chirp signals whose frequency varies with time to modulate information, which achieves better performance than the OFDM technique under doubly dispersive channels. However, the two-dimensional representation of the DD channel in OTFS incurs significant pilot overhead, and the diversity gain that OCDM can obtain depends on specific channel profiles.
Against this background, affine frequency division multiplexing (AFDM) has been proposed based on the discrete affine Fourier transform (DAFT) [16], which can also combat the time-frequency doubly selective fading, and more importantly has less complexity to implement than OTFS since it requires only one-dimension transformation. DAFT is defined as a generalized discrete form of the discrete Fourier transform (DFT) with a chirp-like basis specified by dual adjustable parameters, termed pre-chirp and post-chirp parameters, respectively. In AFDM, data symbols are multiplexed onto chirp-like subcarriers through DAFT and inverse DAFT (IDAFT), which can separate the doubly dispersive channel into a sparse, quasi-static channel with a comprehensive DD channel representation by appropriately setting the chirp parameters. Therefore, AFDM achieves similar performance to OTFS, and demonstrates superior performance over OFDM and OCDM under doubly selective channels [17].
There has been preliminary literature on AFDM [18, 19, 20]. A low-complexity embedded pilot-aided diagonal reconstruction (EPA-DR) channel estimation scheme was proposed in [18], which calculates the AFDM effective channel matrix directly without estimating the three channel parameters, eliminating the inherently severe inter-Doppler interference. In [19], the authors investigated the AFDM-empowered sparse code multiple access (SCMA) systems to support massive connectivity in high-mobility environments. An AFDM-based integrated sensing and communications (ISAC) system was studied in [20], demonstrating that the AFDM-ISAC system can maintain excellent sensing performance even under significant Doppler shifts. The existing literature mostly explored the channel estimation, multiple access and ISAC issues under the classical AFDM architecture, whilst studies regarding further optimization/enhancement of the AFDM waveform are still at their infancies.
One promising research direction is to incorporate the index modulation (IM) philosophy for spectral and energy efficiency improvement [21, 22], which conveys energy-free bits through the activation patterns of transmit entities, e.g., subcarriers [23], time slots [24], pulse positions [25], antennas [26], etc. In [27, 28], Tao et al. presented an IM-assisted scheme, which conveys energy-free information bits through the activation patterns of the subsymbols in the DAFT-domain, verifying that index bits have stronger diversity protection than modulation bits. A multicarrier system using the activation patterns of AFDM chirp subcarriers as indices was developed in [29], which indicates the potential of IM-assisted AFDM technology in enhancing bit error rate (BER) and energy efficiency performance. However, existing research has concentrated on post-chirp parameters in AFDM, with less attention paid to the considerable flexibility and degrees of freedom (DoFs) that the pre-chirp parameter offers.
Distinctively, this paper proposes a novel AFDM scheme with the pre-chirp-domain index modulation (AFDM-PIM) to enhance both spectral and energy efficiencies. Furthermore, performance evaluation of the proposed AFDM-PIM structure, including pairwise error probability (PEP) analysis and diversity analysis, is performed, and the numerical selection of the pre-chirp parameters is analyzed and optimized. The main contributions of this work are highlighted as follows:
  • We prove that the subcarrier orthogonality is maintained by applying distinct pre-chirp parameters to different subcarriers during the AFDM modulation process. Based on this property, each AFDM subcarrier is constructed with a unique pre-chirp signal corresponding to the incoming bits. This configuration allows for the embedding of additional bits into the index patterns of pre-chirp parameter assignment without additional energy consumption.
  • We derive the input-output relationship of the proposed AFDM-PIM scheme in the DAFT domain, and the asymptotically tight upper bounds on the average bit error probability (BEP) with the maximum likelihood (ML) detection based on the PEP analysis. Furthermore, we validate that the proposed AFDM-PIM scheme can achieve full diversity order under doubly dispersive channels.
  • We propose an optimal pre-chirp alphabet design to enhance the BER performance via particle swarm optimization (PSO) algorithm. It is verified via extensive simulations that the optimized pre-chirp parameter alphabet results in a much better BER performance than the heuristic selection of pre-chirp parameter values. Our results also demonstrate that the proposed AFDM-PIM scheme is superior to classical AFDM and IM-aided OFDM algorithms in terms of BER performance.
The rest of the paper is organized as follows. In Section II, the AFDM system model is introduced. Section III details the proposed AFDM-PIM scheme. The performance analysis of AFDM-PIM under doubly dispersive channels is presented in Section IV, which includes the PEP and diversity analysis. The pre-chirp parameter optimization is provided in Section V. The simulation results and discussions are offered in Section VI, and Section VII draws the conclusions.
[…]
Figure 1: The block diagram of AFDM system.
II AFDM System Model
The general system model of AFDM is presented in Fig. 1. For clarity, we provide a concise review of the fundamental concept of AFDM [17]. The transmitted bit stream is initially mapped onto a symbol vector, denoted as $\mathbf{x}_{A}\!=\!\left[x_{A}[0],x_{A}[1],\ldots,x_{A}[N-1]\right]^{\mathrm{T}}\!\in\!\mathbb{C}^{N\times 1}$, comprising $N$ $M$-ary phase shift keying (PSK) symbols in the DAFT domain. The resulting signals are then converted to time-domain representations with an $N$-point IDAFT, formulated as
$s_{A}[n]=\frac{1}{\sqrt{N}}\sum_{m=0}^{N-1}x_{A}[m]e^{\textsf{j}2\pi\left(c_{1}n^{2}+c_{2}m^{2}+\frac{mn}{N}\right)},$ (1)
where $s_{A}[n]$ is the time domain signal, and $m,n\!\in\!\{0,1,\ldots,N-1\}$, while $c_{1}$ and $c_{2}$ are the post-chirp and pre-chirp parameters of the DAFT, respectively.
Similarly to OFDM, AFDM also necessitates the insertion of prefix to address the multi-path problem. By leveraging the inherent periodicity characteristic of DAFT, a chirp-periodic prefix (CPP) is inserted to serve the function analogous to the cyclic prefix (CP) in OFDM, which is given by
$s_{A}[n]\!=\!s_{A}[N\!+\!n]e^{-\textsf{j}2\pi c_{1}(N^{2}+2Nn)},\,n\!=\!-L_{\mathrm{cp}},\ldots,-1,\!$ (2)
where $L_{\mathrm{cp}}$ is the length of the CPP.
Under high-mobility scenarios, the transmitted signals may experience time-frequency doubly dispersive channel attributed to the severe Doppler shift and multi-path effects, which can be modeled as
$h(\tau,\nu)=\sum_{p=1}^{P}h_{p}\delta(\tau-\tau_{p})\delta(\nu-\nu_{p}),$ (3)
where $P$ is the number of the paths, $\nu_{p}$ and $\tau_{p}$ are the Doppler shift and delay of the $p$-th path, respectively, while $h_{p}\!\sim\!\mathcal{CN}(0,1/P)$ is the $p$-th path’s channel coefficient. The normalized delay and Doppler shift are given by $d_{p}\!=\!\tau_{p}\Delta f$ and $\alpha_{p}\!=\!NT\nu_{p}$, respectively, where $\Delta f$ is the AFDM subcarrier spacing and $T$ is the sampling interval with $T\Delta f\!=\!1$. Furthermore, $\alpha_{p}\!\in\![-\alpha_{\max},\,\alpha_{\max}]$ and $d_{p}\!\in\![0,\,d_{\max}]$, where $\alpha_{\max}$ and $d_{\max}$ denote the maximum Doppler shift and maximum delay, respectively [27]. For simplicity and without loss of generality, we mainly consider integer values of $\alpha_{p}$ in this paper.
At the receiver, by discarding the CPP, the received time domain symbols can be expressed as
$r_{A}[n]=\sum_{p=1}^{P}h_{p}s_{A}[n-d_{p}]e^{-\textsf{j}2\pi\nu_{p}n}+w[{n}],$ (4)
where $w[{n}]\!\sim\!\mathcal{CN}\left(0,N_{0}\right)$ is the additive white Gaussian noise (AWGN). After the $N$-point DAFT, the received AFDM signals in the DAFT domain can be expressed as
$y_{A}[\bar{m}]=\frac{1}{\sqrt{N}}$ $\sum_{n=0}^{N-1}r_{A}[n]e^{-\textsf{j}2\pi\left(c_{1}n^{2}+c_{2}\bar{m}^{2}+n\bar{m}/N\right)}$
$\bar{m}=0,1,\ldots,N-1,$ (5)
where $\bar{m}$ is the indices in the DAFT domain. In matrix form, the received AFDM signals can be further written as
$\mathbf{y}_{A}=\sum_{p=1}^{P}h_{p}\mathbf{A}\Gamma_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\bm{\Pi}^{d_{p}}\mathbf{A}^{\mathrm{H}}\mathbf{x}_{A}+\mathbf{w},$ (6)
where the diagonal matrix $,\bm{\Delta}_{\nu_{p}}\!=\!\operatorname{diag}\big{(}e^{-\textsf{j}2\pi\nu_{p}\cdot 0},e^{-\textsf{j}2\pi\nu_{p}\cdot 1},$$\ldots,e^{-\textsf{j}2\pi\nu_{p}\cdot(N-1)}\big{)}$ represents the Doppler effect, $\mathbf{A}$ is the DAFT matrix, and $\mathbf{w}$ is the noise vector, while $\bm{\Pi}$ represents the forward cyclic-shift matrix given by
$\mathbf{\Pi}=\left[\begin{array}[]{cccc}0&\cdots&0&1\\ 1&\cdots&0&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&1&0\end{array}\right]_{N\times N},$ (7)
and $\Gamma_{\text{CPP}_{p}}=\operatorname{diag}\big{(}\omega_{p,0},\omega_{p,1},\ldots,\omega_{p,N}\big{)}$ is the $N\times N$ diagonal matrix for CPP with
$\omega_{p,n}=\begin{cases}e^{-\textsf{j}2\pi c_{1}(N^{2}-2N(d_{p}-n))},&n<d_{p},\\ \qquad\qquad 1,&n\geq d_{p}.\end{cases}$ (8)
Upon receiving the signal $\mathbf{y}_{A}$, the ML detector can be employed for signal detection.
III Proposed AFDM-PIM Scheme
We first analyze the orthogonality of AFDM subcarriers and then derive the proposed AFDM-PIM framework. The input-output relation and parameter settings of the AFDM-PIM are also presented.
III-A Orthogonality Analysis of AFDM Subcarriers
Following the modulation process of AFDM, (1) can also be expressed as
$s_{A}[n]=\sum_{m=0}^{N-1}x_{A}[m]\phi_{n}(m),~{}n=0,1,\ldots,N-1,$ (9)
where $\phi_{n}(m)$ denotes the $m$-th chirp-like subcarrier given by
$\phi_{n}(m)=\frac{1}{\sqrt{N}}\,e^{\textsf{j}2\pi\left(c_{1}n^{2}+c_{2}m^{2}+\frac{mn}{N}\right)}.$ (10)
The following Theorem demonstrates the flexibility of the pre-chirp parameter $c_{2}$ assignment for different subcarriers.
Theorem 1.
Applying distinct $c_{2}$ to different subcarriers in the AFDM modulation process will still preserve their orthogonality.
Proof.
The inner product between two subcarriers of AFDM, which utilize the same post-chirp parameter $c_{1}$ but distinct values of pre-chirp parameter $c_{2}$, designated as $\phi_{n}^{c_{1},c_{2,1}}(m)$ and $\phi_{n}^{c_{1},c_{2,2}}(m)$, respectively, is given by
$\sum_{n=0}^{N-1}\phi_{n}^{c_{1},c_{2,1}}(m_{1})\big{(}\phi_{n}^{c_{1},c_{2,2}}(m_{2})\big{)}^{*}$
$\hskip 11.38109pt=\frac{1}{N}e^{-\textsf{j}2\pi\left(c_{2,1}m_{1}^{2}-c_{2,2}m_{2}^{2}\right)}\sum_{n=0}^{N-1}e^{-\textsf{j}\frac{2\pi}{N}(m_{1}-m_{2})n}$
$\hskip 11.38109pt=\frac{1}{N}e^{-\textsf{j}2\pi\left(c_{2,1}m_{1}^{2}-c_{2,2}m_{2}^{2}\right)}\frac{1-e^{-\textsf{j}2\pi N\left(\frac{m_{1}-m_{2}}{N}\right)}}{1-e^{-\textsf{j}2\pi\left(\frac{m_{1}-m_{2}}{N}\right)}}$
=$\hskip 11.38109pt=\begin{cases}1,&m_{1}=m_{2},\\ 0,&\text{otherwise.}\end{cases}$ (11)
It is evident that the orthogonality among AFDM subcarriers is maintained when different values of $c_{2}$ are employed. ∎
Inspired by Theorem 1, we design the AFDM-PIM scheme, as shown in Fig. 2, which utilizes the flexibility of $c_{2}$ assignment to convey additional information bits.
Figure 2: Transceiver structure of the proposed AFDM-PIM scheme.
TABLE I: Mapping rule between the index bits and the PCPs in the Case of $N_{c}=4$ and $\lambda=4$.
Index bits PCPs for Each Group
subcarrier 1 subcarrier 2 subcarrier 3 subcarrier 4
$0$ $c_{2}^{(1)}$ $c_{2}^{(2)}$ $c_{2}^{(3)}$ $c_{2}^{(4)}$
0001 $c_{2}^{(1)}$ $c_{2}^{(2)}$ $c_{2}^{(4)}$ $c_{2}^{(3)}$
0010 $c_{2}^{(1)}$ $c_{2}^{(3)}$ $c_{2}^{(2)}$ $c_{2}^{(4)}$
0011 $c_{2}^{(1)}$ $c_{2}^{(3)}$ $c_{2}^{(4)}$ $c_{2}^{(2)}$
0100 $c_{2}^{(1)}$ $c_{2}^{(4)}$ $c_{2}^{(2)}$ $c_{2}^{(3)}$
0101 $c_{2}^{(1)}$ $c_{2}^{(4)}$ $c_{2}^{(3)}$ $c_{2}^{(2)}$
0110 $c_{2}^{(2)}$ $c_{2}^{(1)}$ $c_{2}^{(3)}$ $c_{2}^{(4)}$
0111 $c_{2}^{(2)}$ $c_{2}^{(1)}$ $c_{2}^{(4)}$ $c_{2}^{(3)}$
1000 $c_{2}^{(2)}$ $c_{2}^{(3)}$ $c_{2}^{(1)}$ $c_{2}^{(4)}$
1001 $c_{2}^{(2)}$ $c_{2}^{(3)}$ $c_{2}^{(4)}$ $c_{2}^{(1)}$
1010 $c_{2}^{(2)}$ $c_{2}^{(4)}$ $c_{2}^{(1)}$ $c_{2}^{(3)}$
1011 $c_{2}^{(2)}$ $c_{2}^{(4)}$ $c_{2}^{(3)}$ $c_{2}^{(1)}$
1100 $c_{2}^{(3)}$ $c_{2}^{(1)}$ $c_{2}^{(2)}$ $c_{2}^{(4)}$
1101 $c_{2}^{(3)}$ $c_{2}^{(1)}$ $c_{2}^{(4)}$ $c_{2}^{(2)}$
1110 $c_{2}^{(3)}$ $c_{2}^{(2)}$ $c_{2}^{(1)}$ $c_{2}^{(4)}$
$1$ $c_{2}^{(3)}$ $c_{2}^{(2)}$ $c_{2}^{(4)}$ $c_{2}^{(1)}$
III-B Transmitter
Consider the same AFDM symbol vector comprising $N$ $M$-ary constellation symbols in the DAFT domain as in Section II. At the transmitter, unlike classical AFDM, the $N$ AFDM subcarriers are divided into $G$ groups, with each group comprising $N_{c}\!=\!N/G$ chirp subcarriers. The total $B$ information bits are split into $G$ parallel streams of $b\!=\!B/G$ bits for each subcarrier group. Each $b$-bit stream is further segmented into $b_{1}$ symbol bits and $b_{2}$ index bits, i.e., $b\!=\!b_{1}+b_{2}$. Within the $g$-th group ($1\leq g\leq G$), the $b_{1}\!=\!N_{c}\log_{2}(M)$ symbol bits are conveyed by $N_{c}$ $M$-ary symbols, denoted as $\mathbf{x}^{g}\!=\![x^{g}[0],x^{g}[1],\ldots,x^{g}[N_{c}-1]]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$. On the other hand, each subcarrier is assigned with a unique $c_{2}$ value from a finite alphabet of $\lambda$ legitimate $c_{2}$ realizations, i.e., $\mathcal{P}_{c}\!=\!\left\{{c_{2}^{(1)},c_{2}^{(2)},\ldots,c_{2}^{(\lambda)}}\right\}$. Specifically, the pre-chirping pattern (PCP) of the $c_{2}$ values in the $g$-th group, denoted by $\mathrm{P}_{c_{2}}^{g}\!=\![c_{2,N_{c}\left(g-1\right)},c_{2,N_{c}\left(g-1\right)+1},\ldots,c_{2,N_{c}g-1}]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$ is determined by the $b_{2}$ index bits according to the pre-defined relationship between the index-bit stream and the permutations of $N_{c}$ elements from $\mathcal{P}_{c}$, where $c_{2,m}\!\in\!\mathcal{P}_{c}$ represents the pre-chirp parameter of the $m$-th subcarrier. $\mathrm{P}_{c_{2}}\!=\!\left[\mathrm{P}_{c_{2}}^{1},\mathrm{P}_{c_{2}}^{2},\ldots,\mathrm{P}_{c_{2}}^{G}\right]$ represents the PCP for all the groups (PCPG), and all the possible PCPGs are combined into a set, termed $\mathcal{S}_{p}$. Table I exemplifies the mapping rule in the case of $N_{c}=4$ and $\lambda=4$. Every $b_{2}$ index bits correspond to one row in Table I. Hence, aside from the classical PSK modulation, additional $b_{2}$ information bits can be implicitly conveyed by the indices of $\mathrm{P}_{c_{2}}^{g}$, and
$b_{2}=\begin{cases}\left\lfloor\log_{2}(\mathcal{C}_{\lambda,N_{c}}N_{c}!)\right\rfloor,\vspace{0.1cm}&\lambda\geq N_{c},\\ \left\lfloor\log_{2}(\lambda!)\right\rfloor\frac{N_{c}}{\lambda},&\lambda<N_{c}\text{ \& }\lambda\mid N_{c},\\ \left\lfloor\log_{2}(\mathcal{C}_{\lambda,N_{c}}\lambda!\lambda^{(N_{c}-\lambda)})\right\rfloor,&\text{otherwise,}\end{cases}$ (12)
where $\mathcal{C}_{\lambda,N_{c}}$ is defined as
$\mathcal{C}_{\lambda,N_{c}}=\binom{\max\left(\lambda,N_{c}\right)}{\min\left(\lambda,N_{c}\right)}.$ (13)
After the mapping for all the groups, the time-domain transmitted signals can be generated through $N$-point IDAFT operation, expressed as
$s[n]=\frac{1}{\sqrt{N}}\sum_{m=0}^{N-1}x[m]e^{\textsf{j}2\pi\left(c_{1}n^{2}+c_{2,m}m^{2}+nm/N\right)}.$ (14)
The matrix form of (14) can be formulated as
$\mathbf{s}=\mathbf{A}^{\mathrm{H}}\mathbf{x}=\mathbf{\Lambda}_{c_{1}}^{\mathrm{H}}\mathbf{F}^{\mathrm{H}}\bm{\Lambda}_{c_{2}}^{\mathrm{H}}\mathbf{x},$ (15)
where $\mathbf{x}\!=\!\left[\big{(}\mathbf{x}^{1}\big{)}^{\rm T},\big{(}\mathbf{x}^{2}\big{)}^{\rm T},\ldots,\big{(}\mathbf{x}^{G}\big{)}^{\rm T}\right]^{\mathrm{T}}$, and $\bm{\Lambda}_{c_{2}}$ and $\bm{\Lambda}_{c_{1}}$ represent the pre-chirp and post-chirp diagonal matrices, respectively, expressed as
$\bm{\Lambda}_{c_{2}}=\operatorname{diag}\left(e^{-\textsf{j}2\pi c_{2,n}n^{2}},n=0,1,\ldots,N-1\right),$ (16)
$\bm{\Lambda}_{c_{1}}=\operatorname{diag}\left(e^{-\textsf{j}2\pi c_{1}n^{2}},n=0,1,\ldots,N-1\right),$ (17)
while $\mathbf{F}$ denotes the DFT matrix with elements $\mathbf{F}(m,n)=e^{-\textsf{j}2\pi mn/N}/\sqrt{N}$, $m,n=0,1,\ldots,N-1$.
Like classical AFDM, our proposed AFDM-PIM also requires the CPP to mitigate the effects of multi-path propagation effectively. Without loss of generality, the length of the CPP is assumed to be greater than the maximum channel delay spread.
III-C Receiver
The received time-domain signals after removing the CPP can be written as
$r[n]=\sum_{p=1}^{P}h_{p}s[n-d_{p}]e^{-\textsf{j}2\pi\nu_{p}n}+w_{r}[n],$ (18)
where $w_{r}[n]\sim\mathcal{CN}(0,N_{0})$ is the AWGN. The matrix form of (18) is given by
$\mathbf{r}=\mathbf{H}\mathbf{s}+\mathbf{w}=\sum_{p=1}^{P}h_{p}\Gamma_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\mathbf{\Pi}^{d_{p}}\mathbf{s}+\mathbf{w}_{r},$ (19)
where $\mathbf{w}_{r}\!=\![w_{r}[0],w_{r}[1],\ldots,w_{r}[N-1]]^{\mathrm{T}}$ is the time-domain noise vector. As given in Section II, $\bm{\Delta}_{\nu_{p}}$ represents the Doppler effect, $\mathbf{\Pi}$ is the forward cyclic-shift matrix with $\mathbf{\Pi}^{d_{p}}$ modeling the delay extension, and $\Gamma_{\mathrm{CPP}_{p}}$ is the effective CPP matrix.
By applying the DAFT operation, the received DAFT-domain symbols are obtained as
$y[\bar{m}]=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}r[n]e^{-\textsf{j}2\pi\left(c_{1}n^{2}+c_{2,\bar{m}}\bar{m}^{2}+n\bar{m}/N\right)},$ (20)
which can also be grouped into a vector as
$\mathbf{y}=$ $\mathbf{A}\mathbf{r}=\sum_{p=1}^{P}h_{p}\mathbf{A}\mathbf{\Gamma}_{\mathrm{CPP}_{p}}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{d_{p}}\mathbf{A}^{\mathrm{H}}\mathbf{x}+\mathbf{A}\mathbf{w}_{r}$
$=$ $\mathbf{H}_{\textnormal{eff}}\mathbf{x}+\mathbf{w},$ (21)
where $\mathbf{H}_{\textnormal{eff}}$ is the effective channel matrix in the DAFT-domain and $\mathbf{w}=\mathbf{A}\mathbf{w}_{r}$.
The time-frequency doubly dispersive channel can be estimated through pilot-aided channel estimation algorithms [18, 30, 31]. This paper will not provide further elaboration on this subject for brevity. Given the estimated $\mathbf{\hat{H}}_{\mathrm{eff}}$ for the effective channel matrix $\mathbf{H}_{\textnormal{eff}}$, the ML data detection is formulated as the following optimization
$\left(\hat{\mathbf{x}},\hat{\mathrm{P}}_{c_{2}}\right)=\arg\min\limits_{\forall\mathbf{x},\mathrm{P}_{c_{2}}}\left\|\mathbf{y}-\mathbf{\hat{H}}_{\mathrm{eff}}\mathbf{x}\right\|^{2}.$ (22)
III-D Input-Output Relation and Parameter Settings
Substituting (14) and (18) into (20), the input-output relation of AFDM-PIM can be obtained as
$y[\bar{m}]\!=\!\frac{1}{N}\sum_{p=1}^{P}\!\sum_{m=0}^{N-1}\!h_{p}\xi_{(p,\bar{m},m)}\eta_{(p,\bar{m},m)}x[m]+w[\bar{m}],$ (23)
where
$\xi_{(p,\bar{m},m)}=e^{\textsf{j}\frac{2\pi}{N}\left(Nc_{2,m}m^{2}-Nc_{2,\bar{m}}\bar{m}^{2}-md_{p}+Nc_{1}d_{p}^{2}\right)},$ (24)
$\eta_{(p,\bar{m},m)}=$ $\sum_{n=0}^{N-1}e^{-\textsf{j}\frac{2\pi}{N}((\bar{m}-m+\alpha_{p}+2Nc_{1}d_{p})n)}$
$=$ $\frac{e^{-\textsf{j}2\pi(\bar{m}-m+\alpha_{p}+2Nc_{1}d_{p})}-1}{e^{-\textsf{j}\frac{2\pi}{N}(\bar{m}-m+\alpha_{p}+2Nc_{1}d_{p})}-1}.$ (25)
In matrix representation, (23) can be rewritten as
$\mathbf{y}=\sum_{p=1}^{P}h_{p}\mathbf{H}_{p}\mathbf{x}+\mathbf{w},$ (26)
where the elements of $\mathbf{H}_{p}$ are given by
$H_{p}[\bar{m},m]=$ $\frac{1}{N}\xi_{(p,\bar{m},m)}\eta_{(p,\bar{m},m)}$
$=$ $\begin{cases}\xi_{(p,\bar{m},m)},&m=(\bar{m}+\mathrm{loc}_{p})_{N},\\ \qquad 0,&\text{otherwise,}\end{cases}$ (27)
where $\mathrm{loc}_{p}=(\alpha_{p}+2Nc_{1}d_{p})_{N}$. Since the range of $\mathrm{loc}_{p}$ is $[-\alpha_{\max}+2Nc_{1}d_{p},\,\alpha_{\max}+2Nc_{1}d_{p}]$, we define $\mathrm{loc}_{p}\in\mathbb{K}_{p}$, with $\mathbb{K}_{p}=\{-\alpha_{\mathrm{max}}+2Nc_{1}d_{p},\ldots,\alpha_{\mathrm{max}}+2Nc_{1}d_{p}\}$.
It can be seen from (III-D) that for the two chirp parameters of AFDM-PIM, only the post-chirp parameter exerts an influence on the positions of non-zero entries in the matrix $\mathbf{H}_{p}$ determined by $\mathrm{loc}_{p}$, which is independent of the pre-chirp parameter. Therefore, like classical AFDM, full diversity order can be obtained by the proposed AFDM-PIM under doubly dispersive channels by adjusting $c_{1}$ to avoid possible overlap between non-zero elements of $\mathbf{H}_{i}$ and $\mathbf{H}_{j}$ $(i\neq j)$. Specifically, this requires that the intersection between the corresponding ranges for $\mathrm{loc}_{i}$ and $\mathrm{loc}_{j}$ is empty, i.e.,
$\mathbb{K}_{i}\cap\mathbb{K}_{j}=\emptyset.$ (28)
Without loss of generality, assume that $d_{i}\leq d_{j}$. Then the constraint (28) can be transformed into
$c_{1}>\frac{2\alpha_{\max}}{2N(d_{j}-d_{i})}.$ (29)
Since the minimum value of $(d_{j}-d_{i})$ with $i\neq j$ equals one, it can be concluded that $c_{1}$ can be set as:
$c_{1}=\frac{2\alpha_{\max}+1}{2N}.$ (30)
Following the configuration (30) and $(d_{\max}+1)(2\alpha_{\max}+1)\leq N$, the channel paths with different delays or Doppler shifts can be distinguished within the DAFT domain, as illustrated in Fig. 3, which shows an example of the effective channel matrix of an AFDM-PIM system under a three-path channel. The rigorous diversity analysis will be provided in Section IV.
Figure 3: Example of the effective channel matrix of a three-path channel.
IV Performance Analysis
We first derive the average BEP (ABEP) upper bounds for our AFDM-PIM scheme with the ML detection. Then we analyze the diversity order achieved by the system.
IV-A Error Performance Analysis
To facilitate the analysis, the received DAFT-domain signal (26) can be rewritten as
$\mathbf{y}=\mathbf{\Phi}(\mathbf{x})\mathbf{h}+\mathbf{w},$ (31)
where $\mathbf{\Phi}(\mathbf{x})\!=\![\mathbf{H}_{1}\mathbf{x},\mathbf{H}_{2}\mathbf{x},\ldots,\mathbf{H}_{P}\mathbf{x}]\!\in\!\mathbb{C}^{N\times P}$ and $\mathbf{h}\!=\![h_{1},h_{2},\ldots,h_{P}]^{\rm T}\!\in\!\mathbb{C}^{P\times 1}$. The conditional PEP (CPEP) between $\mathbf{\Phi}(\mathbf{x})$ and its estimate $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ can be calculated as
$\Pr([\mathbf{x},\mathbf{\Phi}]\!\to\![\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]|\mathbf{h})=\Pr(\|\mathbf{y}\!-\!\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\mathbf{h}\|^{2}\!<\!\|\mathbf{y}\!-\!\mathbf{\Phi}(\mathbf{x})\mathbf{h}\|^{2})$
$\hskip 28.45274pt=\Pr\left(\mathbf{\chi}>\big{\|}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\big{\|}^{2}\right),$ (32)
where $\chi\!=\!\mathbf{w}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\!+\!\mathbf{h}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}^{\rm H}\mathbf{w}$. Since $\chi$ follows Gaussian distribution with variance $2N_{0}\big{\|}(\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x}))\mathbf{h}\big{\|}^{2}$, the CPEP can be expressed as
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]|\mathbf{h}\big{)}=Q\biggl{(}\sqrt{\frac{\delta}{2N_{0}}}\biggr{)},$ (33)
where $\delta\!=\!\left\|\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\right\|^{2}\!=\!\mathbf{h}^{\mathrm{H}}\mathbf{\Psi}\mathbf{h}$ with $\mathbf{\Psi}\!=\!\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}^{\mathrm{H}}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}$. According to [32], $Q$-function can be approximated as
$Q(x)\approx\frac{1}{12}e^{-x^{2}/2}+\frac{1}{4}e^{-2x^{2}/3}.$ (34)
Hence the CPEP can be approximated as
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]|\mathbf{h}\big{)}\approx\frac{1}{12}e^{-\varsigma_{1}\delta}+\frac{1}{4}e^{-\varsigma_{2}\delta},$ (35)
with $\varsigma_{1}\!=\!\frac{1}{4N_{0}}$ and $\varsigma_{2}\!=\!\frac{1}{3N_{0}}$. Consequently, the unconditional PEP (UPEP) can be calculated as
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]$ $\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}=\mathrm{E}\Big{(}\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]|\mathbf{h}\big{)}\Big{)}$
$\approx\int_{0}^{+\infty}\biggl{(}\frac{1}{12}e^{-\varsigma_{1}\delta}+\frac{1}{4}e^{-\varsigma_{2}\delta}\biggr{)}p_{\delta}(\delta)\mathrm{d}\delta,$ (36)
where $p_{\delta}(\delta)$ denotes the probability density function (PDF) of $\delta$. Noting $p_{\delta}(\delta)=0$ for $\delta<0$ and utilizing the definition of the moment-generating function (MGF) $M_{\eta}(s)\!=\!\mathrm{E}\big{(}e^{s\eta}\big{)}\!=\!\int_{-\infty}^{+\infty}e^{s\eta}p_{\eta}(\eta)\mathrm{d}\eta$, we can calculate the UPEP (IV-A) as
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}\approx\frac{1}{12}M_{\delta}(-\varsigma_{1})+\frac{1}{4}M_{\delta}(-\varsigma_{2}).$ (37)
Theorem 2 ([33]).
For an $N\times N$ Hermitian matrix $\mathbf{Q}$ and an $N\times 1$ zero-mean complex-valued random vector $\mathbf{v}$ with covariance matrix $\mathbf{L}$, the characteristic function of the quadratic form $f\!=\!\mathbf{v}^{\mathrm{H}}\mathbf{Q}\mathbf{v}$ can be expressed as
$\varphi_{f}(t)=|\mathbf{I}-\textsf{j}\cdot t\mathbf{L}\mathbf{Q}|^{-1}=\prod_{\iota=1}^{\kappa}\frac{1}{1-\textsf{j}\cdot t\lambda_{\iota}^{(\mathbf{L}\mathbf{Q})}},$ (38)
where $\lambda_{\iota}^{(\mathbf{L}\mathbf{Q})}$ is the $\iota$-th non-zero eigenvalue of matrix $\mathbf{L}\mathbf{Q}$, and $\kappa$ denotes the number of the non-zero eigenvalues, i.e., the rank of $\mathbf{L}\mathbf{Q}$. Using the relationship $M_{f}(s)=\varphi_{f}(-\textsf{j}\cdot s)$, the MGF of $f$ is given by
$M_{f}(s)=\prod_{\iota=1}^{\kappa}\frac{1}{1-s\lambda_{\iota}^{(\mathbf{L}\mathbf{Q})}}.$ (39)
Since $\mathbf{h}$ is a zero mean complex vector with the covariance matrix $\frac{1}{P}\mathbf{I}_{P}$ and $\mathbf{\Psi}$ is a Hermitian matrix, according to Theorem 2, the UPEP (37) can be further expressed as
$\Pr\!\big{(}[\mathbf{x},\mathbf{\Phi}]\!\to\!\big{[}\hat{\mathbf{x}},\hat{\mathbf{\Phi}}\big{]}\big{)}\!\approx\!\frac{1}{12}\prod_{\iota=1}^{\kappa}\frac{1}{1+\frac{\lambda_{\iota}^{(\mathbf{\Psi})}}{4PN_{0}}}\!+\!\frac{1}{4}\prod_{\iota=1}^{\kappa}\frac{1}{1\!+\!\frac{\lambda_{\iota}^{(\mathbf{\Psi})}}{3PN_{0}}},$ (40)
where we assume that the rank of $\mathbf{\Psi}$ is $\kappa$.
Moreover, based on the UPEP (40), the ABEP upper bound for the proposed AFDM-PIM scheme can be calculated by
[…]
×$\hskip 71.13188pt\times\tau\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)},$ (41)
where $\tau\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}$ represents the number of error bits caused by the corresponding pairwise error event.
IV-B Diversity Analysis
At the high signal-to-noise ratio (SNR) region, the approximation of (40) can be simplified as
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}$
$\hskip 14.22636pt\approx\left(\prod_{\iota=1}^{\kappa}\lambda_{\iota}^{(\mathbf{\Psi})}\right)^{-1}\left(\frac{(4P)^{\kappa}}{12}+\frac{(3P)^{\kappa}}{4}\right)\mathrm{SNR}^{-\kappa},$ (42)
where $\mathrm{SNR}$ is defined as $1/N_{0}$. As a result, the diversity order $\mu$ of the proposed AFDM-PIM scheme equals to the minimum value of $\kappa$, i.e.,
$\mu=\min\,\mathrm{rank}(\mathbf{\Psi}).$ (43)
Since the eigenvalues of Hermitian matrix $\mathbf{\Psi}$ can be calculated as the square of the singular values of $\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\big{)}$, $\mathrm{rank}(\mathbf{\Psi})\!=\!\mathrm{rank}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\big{)}$ and the diversity order $\mu$ can also be expressed as
$\mu=\min\,\mathrm{rank}\left(\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\right).$ (44)
Defining $\mathbf{\Phi}\big{(}\bar{\delta}\big{)}\!=\!\hat{\mathbf{\Phi}}\big{(}\hat{\mathbf{x}}\big{)}-\mathbf{\Phi}(\mathbf{x})$, the full diversity order analysis for the AFDM-PIM can be transformed into the full rank analysis of $\mathbf{\Phi}\big{(}\bar{\delta}\big{)}$ as stated in Theorem 3 below.
Theorem 3.
The proposed AFDM-PIM scheme is capable of achieving the full diversity order if the following two conditions are met.
Condition 1: The number of the paths satisfies
$P\leq(d_{\mathrm{max}}+1)(2\alpha_{\mathrm{max}}+1)\leq N.$ (45)
Condition 2: The pre-chirp parameters in $\mathcal{P}_{c}$ take the irrational numbers.
Proof.
See Appendix -A. ∎
V Parameter Optimization
In this section, we establish an optimization problem for the optimal $c_{2}$ alphabet design, where the PSO algorithm is employed, to enhance the BER performance.
V-A Problem Formulation
For tractable analysis, the pre-chirp alphabet design is investigated based on the optimal BER detector [34], i.e., the ML detector based on the signal model (31), expressed as
$\big{(}\hat{\mathbf{x}},\hat{\mathrm{P}}_{c_{2}}\big{)}=\arg\min_{\forall\mathbf{x},\mathrm{P}_{c_{2}}}\|\mathbf{y}-\mathbf{\Phi}(\mathbf{x})\mathbf{h}\|^{2}.$ (46)
To achieve the optimal $c_{2}$ alphabet, the minimum Euclidean distance (MED) between different realizations of $\mathbf{\Phi}(\mathbf{x})$ should be maximized. The Euclidean distance between two realizations of $\mathbf{\Phi}(\mathbf{x})$ is formulated as
$O_{k,j}(\mathcal{P}_{c})=\sum_{\mathbf{x}^{\prime}\!,\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})-\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2},$ (47)
where $j$ and $k$ represent the indices of $\mathcal{S}_{p}$, and the corresponding PCPGs are expressed as $\mathrm{P}_{c_{2}}^{(j)}\!=\![c_{2,1},c_{2,2},\ldots,c_{2,N}]$ and $\mathrm{P}_{c_{2}}^{(k)}\!=\![c_{2,1}^{\prime},c_{2,2}^{\prime},\ldots,c_{2,N}^{\prime}]$, respectively, while $r$ is the index of delay and Doppler selection. Specifically, each $r$ corresponds to a specific combination of $P$ paths with different delays and Doppler shifts under doubly dispersive channels with a maximum delay $d_{\text{max}}$ and a normalized Doppler shift $\alpha_{\text{max}}$, and the maximum value of $r$ is $\mathcal{R}=\binom{P_{\mathrm{max}}}{P}$, where $P_{\mathrm{max}}\!=\!(d_{\mathrm{max}}+1)(2\alpha_{\mathrm{max}}+1)$.
According to Theorem 3 and the periodicity brought by $2\pi$ in (16), $c_{2}$ is an irrational number with a principal value range of [0,  1]. To enhance the BER performance of the proposed AFDM-PIM, we formulate the following problem to maximize the MED with optimal $c_{2}$ alphabet design:
$\max_{\mathcal{P}_{c}}\min_{k,j\in[1,2^{b_{2}}]}O_{k,j}(\mathcal{P}_{c}),$ (48a)
[…]
$\hskip 17.07182pt\quad c_{2}\in[0,1],$ (48c)
$\hskip 17.07182pt\quad c_{2}\in\mathbb{I},$ (48d)
where the objective function can be expressed as
$O_{k,j}(\mathcal{P}_{c})\!=\!\sum_{\mathbf{x}^{\prime}\!,\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\!\!\sum_{n=0}^{N-1}\!\left(1\!-\!\cos\left(\psi_{n}+\theta_{n}^{\prime}\!-\!\theta_{n}\right)\right),$ (49)
in which $\psi_{n}$ represents the phase difference between $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$, i.e.,
$e^{\textsf{j}\psi_{n}}=x_{\operatorname{loc}_{p}+n}^{\prime}\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$ (50)
with $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$ denoting the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, respectively, and
$\left\{\begin{array}[]{l}\theta_{n}=2\pi\left(c_{2,\operatorname{loc}_{p}+n}\left(\operatorname{loc}_{p}+n\right)^{2}-c_{2,n}n^{2}\right),\\ \theta_{n}^{\prime}=2\pi\left(c_{2,\operatorname{loc}_{p}+n}^{\prime}\left(\operatorname{loc}_{p}+n\right)^{2}-c_{2,n}^{\prime}n^{2}\right).\end{array}\right.$ (51)
The derivation of $O_{k,j}(\mathcal{P}_{c})$ (49) is given in Appendix -B.
V-B Problem Transformation
V-B1 Transformation of $\mathbf{x}^{\prime}$ and $\mathbf{x}$
Since $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$ take values from the $M$-PSK constellation, $\psi_{n}$ is given by
$\psi_{n}=\frac{2\pi k}{M},k=-(M-1),\ldots,-1,0,1,\ldots,M-1,$ (52)
where each possible value is assigned with equal probability, i.e., $\psi_{n}$ follows a discrete uniform distribution.
For the case where $\mathbf{x}^{\prime}\neq\mathbf{x}$, $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}^{(\mathbf{x}^{\prime}\neq\mathbf{x})}(\mathcal{P}_{c})=\sum_{\mathbf{x}^{\prime}\neq\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(-\cos\left(\psi_{n}+\theta_{n}^{\prime}-\theta_{n}\right)\right)$
$\hskip 8.53581pt=\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\!\!\bigg{(}\!\sum_{k=1-M}^{M-1}\!\!\cos\Big{(}\frac{2\pi k}{M}+\theta_{n}^{\prime}-\theta_{n}\Big{)}\!\bigg{)}=0.$ (53)
Therefore, for the original problem (48), it is sufficient to consider the case of $\mathbf{x}^{\prime}\!=\!\mathbf{x}$. In this case, $\psi_{n}\!=\!0$ for all $n\!=\!0,1,\ldots,N$, and the problem (48) can be expressed as
$\max_{\mathcal{P}_{c}}\min_{k,j\in[1,2^{b_{2}}]}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(1-\cos\left(\theta_{n}^{\prime}-\theta_{n}\right)\right),$ (54a)
[…]
V-B2 Transformation of $\mathrm{P}_{c_{2}}^{(j)}$ and $\mathrm{P}_{c_{2}}^{(k)}$
In the problem (54), each pair of indices of PCPGs, i.e., each pair of $j,k$, corresponds to a specific set of $\mathrm{P}_{c_{2}}^{(j)}$ and $\mathrm{P}_{c_{2}}^{(k)}$. By substituting (51) into the objective function of (54), $O_{k,j}(\mathcal{P}_{c})$ becomes
$O_{k,j}(\mathcal{P}_{c})=\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(1-O^{\prime}\right),$ (55)
where
[…]
It can be observed that in the case of $j\neq k$, i.e., $\mathrm{P}_{c_{2}}^{(j)}\neq\mathrm{P}_{c_{2}}^{(k)}$, as the discrepancy between $\mathrm{P}_{c_{2}}^{(j)}$ and $\mathrm{P}_{c_{2}}^{(k)}$ diminishes, the number of zero values in $O^{\prime}$ increases. Therefore, for the problem (54), it is sufficient to consider the case where $\big{\|}\mathrm{P}_{c_{2}}^{(j)}-\mathrm{P}_{c_{2}}^{(k)}\big{\|}_{0}=2$. Hence, the optimization problem can be further expressed as
$\max_{\mathcal{P}_{c}}\min_{k,j}$ $\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(1-\cos\left(\theta_{n}^{\prime}-\theta_{n}\right)\right),$ (57a)
s.t. $j\neq k,$ (57b)
$\big{\|}\mathrm{P}_{c_{2}}^{(j)}-\mathrm{P}_{c_{2}}^{(k)}\big{\|}_{0}=2,$ (57c)
[…]
$c_{2}\in\mathbb{I}.$ (57e)
V-C Problem Solver
The optimization (57) is a non-convex problem, and it is challenging to obtain a global optimal solution. To this end, the PSO-based algorithm is invoked to obtain a suboptimal solution, attributed to its rapid convergence and exemplary global searching capabilities.
Algorithm 1 PSO-Based Algorithm for $c_{2}$ Alphabet Design
1:$N_{p}$, $\varpi$, $\varrho_{global}$, $\varrho_{local}$, $v_{\max}$, $I_{\max}$, $\lambda$ and all other parameters required to evaluate fitness function;
2:$\mathcal{P}_{c}$;
3:Set $i_{ter}\!=\!0$ and initialize $N_{p}$ particles with positions $\mathcal{P}^{(0)}$ and zero velocities $\mathcal{V}^{(0)}$;
4:Calculate fitness values of all particles by (59), $\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p}}^{(0)}\Big{)}$, $n_{p}=1,2,\ldots,N_{p}$;
5:Initialize local optimal position of each particle $\bm{p}_{n_{p},local}=\bm{p}_{n_{p}}^{(0)}$, calculate global optimal position $\bm{p}_{global}=\arg\max\limits_{1\leq n_{p}\leq N_{p}}\big{(}F_{p}(\bm{p}_{n_{p}}^{(0)})\big{)}$;
6:while $i_{ter}\leq I_{\mathrm{max}}$ do
7:     for $n_{p}=1$ to $N_{p}$ do
8:         Update velocity $\bm{v}_{n_{p}}^{(i_{ter})}$ according to (60);
9:         for $i=1$ to $\lambda$ do
10:              if $\bm{v}_{n_{p}}^{(i_{ter})}[i]>{v}_{\mathrm{max}}$ then
11:                  $\bm{v}_{n_{p}}^{(i_{ter})}[i]\leftarrow{v}_{\max}$;
12:              else if $\bm{v}_{n_{p}}^{(i_{ter})}[i]<-{v}_{\max}$ then
13:                  $\bm{v}_{n_{p}}^{(i_{ter})}[i]\leftarrow-{v}_{\max}$
14:              end if
15:         end for
16:         Update position $\bm{p}_{n_{p}}^{(i_{ter})}$ based on (61);
17:         Calculate fitness value $\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p}}^{(i_{ter})}\Big{)}$ using (59);
18:         if $\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p}}^{(i_{ter})}\Big{)}>\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p},local}\Big{)}$ then
19:              $\bm{p}_{n_{p},local}\leftarrow\bm{p}_{n_{p}}^{(i_{ter})}$;
20:         end if
21:         if $\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p}}^{(i_{ter})}\Big{)}>\mathcal{F}_{P}\Big{(}\bm{p}_{global}\Big{)}$ then
22:              $\bm{p}_{global}\leftarrow\bm{p}_{n_{p}}^{(i_{ter})}$;
23:         end if
24:     end for
25:     $i_{ter}\leftarrow i_{ter}+1$;
26:end while
27:Obtain global optimal position $\mathcal{P}_{c}\leftarrow\bm{p}_{global}$;
28:return $\mathcal{P}_{c}$.
First, a population of $N_{p}$ particles with velocities and positions are initialized. The velocities of the particles are represented by $\mathcal{V}^{(0)}\!=\!\{\bm{v}_{1}^{(0)},\bm{v}_{2}^{(0)}\ldots,\bm{v}_{N_{p}}^{(0)}\}$, which is indicative of the extent of change occurring during the iterative process. The positions of the particles are denoted by $\mathcal{P}^{(0)}\!=\!\{\bm{p}_{1}^{(0)},\bm{p}_{2}^{(0)},\ldots,\bm{p}_{N_{p}}^{(0)}\}$, where each position represents a potential solution for the pre-chirp alphabet, i.e.,
$\bm{p}_{n_{p}}^{(0)}=\mathcal{P}_{c,n_{p}}^{(0)}=\left\{{c_{2,{n_{p}}}^{(0,1)},c_{2,{n_{p}}}^{(0,2)},\ldots,c_{2,{n_{p}}}^{(0,\lambda)}}\right\},$ (58)
in which $n_{p}$ denotes the index of particle. As an initial solution, the first particle $\bm{p}_{1}^{(0)}$ is initialized as a heuristic pre-chirp alphabet, where the elements in $\mathcal{P}_{c}$ are evenly distributed within the interval [0,  1], and the remaining particles are randomly initialized.
Subsequently, the fitness value of each particle is evaluated per the specified utility function. In light of the constraints imposed by (57b) and (57c), a brick wall penalty factor is introduced, and the utility function in the $i_{ter}$-th iteration is defined as
$\mathcal{F}_{P}\Big{(}\bm{p}_{n_{p}}^{(i_{ter})}\Big{)}=\begin{cases}\epsilon,&\text{if}\hskip 5.69046pt\bm{p}_{n_{p}}^{(i_{ter})}\hskip 2.84544pt\text{is feasible,}\\ -1,&\text{otherwise,}\end{cases}$ (59)
where $\epsilon\!=\!\min_{k,j}O_{k,j}(\mathcal{P}_{c})$ and $O_{k,j}(\mathcal{P}_{c})$ is calculated by (55). Then the particle with the greatest fitness value is considered to be the initial global optimal position $\bm{p}_{global}$, and the local optimal position $\bm{p}_{n_{p},local}$ of each particle is initialized as $\bm{p}_{n_{p}}^{(0)}$.
Each particle conveys its local optimal position to other particles during the iteration. The velocity and position of each particle are updated according to
$\bm{v}_{n_{p}}^{(i_{ter})}=$ $\varpi\bm{v}_{n_{p}}^{(i_{ter}-1)}+r_{1}\varrho_{local}\big{(}\bm{p}_{{n_{p}},local}-\bm{p}_{n_{p}}^{(i_{ter}-1)}\big{)}$
$+r_{2}\varrho_{global}\big{(}\bm{p}_{global}-\bm{p}_{n_{p}}^{(i_{ter}-1)}\big{)},$ (60)
$\bm{p}_{n_{p}}^{(i_{ter})}=$ $\bm{p}_{n_{p}}^{(i_{ter}-1)}+\bm{v}_{n_{p}}^{(i_{ter})},$ (61)
respectively, where $\varpi$ is the inertia weight, $\varrho_{global}$ and $\varrho_{local}$ denote the global and local updating coefficients, respectively, while $r_{1}$ and $r_{2}$ are two random values drawn within the interval [0,  1]. The particle velocity is constrained to the range between $-{v}_{\max}$ and ${v}_{\max}$. Afterward, the fitness values of all the particles are evaluated, and the global and local optimal positions, $\bm{p}_{global}$ and $\big{\{}\bm{p}_{n_{p},local}\big{\}}_{n_{p}=1}^{N_{p}}$ , are updated. The above update process is repeated until the maximum number $I_{\mathrm{max}}$ of iterations is reached. The particle with the highest fitness value corresponds to the optimization result $\mathcal{P}_{c}$. The procedure of the proposed PSO-based algorithm is outlined in Algorithm 1.
VI Simulation Results and Analysis
In this section, we carry out simulations to evaluate the performance of the proposed AFDM-PIM schemes, and conduct the BER performance comparison between the proposed AFDM-PIM scheme and other existing counterparts. The accuracy of the PEP-based theoretical derivation is also investigated in comparison with the Monte Carlo simulation results. In simulations, the carrier frequency is set to $f_{c}\!=\!8$ GHz, and the subsymbol spacing in the DAF domain is set to $f_{s}\!=\!1.5$ kHz. Two distinct simulation scenarios are conducted, namely, when the full diversity condition (45) is satisfied and when it is not satisfied. The maximum normalized Doppler shifts for these two scenarios are set to $\alpha_{\mathrm{max}}\!=\!1$ and $\alpha_{\mathrm{max}}\!=\!2$, corresponding to the high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!202.5$ km/h and the ultra high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!405$ km/h, respectively. The Doppler shift of each path is generated according to Jakes Doppler spectrum approach as $\alpha_{p}\!=\!\alpha_{\mathrm{max}}\cos(\theta_{p,d})$, where $\theta_{p,d}\!\in\![-\pi,\,\pi]$ (for integer Doppler cases, the Doppler shift is $\lfloor\alpha_{\mathrm{max}}\cos(\theta_{p,d})\rfloor$). We set the pertinent parameters of Algorithm 1 as ${v}_{\max}\!=\!0.05$, $\varpi\!=\!0.5$, $N_{p}\!=\!200$, $I_{\mathrm{max}}\!=\!300$, $\varrho_{global}\!=\!2$ and $\varrho_{local}\!=\!2$. Unless otherwise specified, the ML detector is employed for both the proposed AFDM-PIM and the classical benchmarks.
Figure 4: BER Performance comparison between the $c_{2}$ alphabet designed by Algorithm 1 and the initial $c_{2}$ solution, given two different numbers of paths.
VI-A Effectiveness of Algorithm 1
First, the effectiveness of the proposed $c_{2}$ alphabet design with Algorithm 1 is validated. Fig. 4 compares the BER performance achieved by the $c_{2}$ alphabet designed by Algorithm 1 and that attained by the initial solution, namely, the first initial particle $\bm{p}_{1}^{(0)}$ for Algorithm 1, given two numbers of paths $P\!=\!2$ and 3. Binary PSK (BPSK) is employed as the modulation scheme for the data bits, and the parameter settings are $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(6,2,3,1,1)$. Note that the full diversity order condition is satisfied, and consequently the BER performance improves with the increase in the number of paths. Fig. 4 indicates that the $c_{2}$ alphabet designed by Algorithm 1 exhibits superior performance by about 3 dB in SNR than the initial solution at the BER level of $10^{-3}$, given $P\!=\!3$. The results of Fig. 4 demonstrate the effectiveness of Algorithm 1.
Figure 5: Performance comparisons between the proposed AFDM-PIM scheme and the classical AFDM scheme given two different spectral efficiencies.
VI-B Performance Comparison with Existing Benchmarks
Fig. 5 compares the BER performance of the proposed AFDM-PIM with that of the classical AFDM scheme [17] under the same doubly dispersive channel with $P\!=\!4$. The proposed AFDM-PIM employs BPSK and quadrature PSK (QPSK) to achieve the spectral efficiency of $2$ and $3$ bit/s/Hz, respectively. The other parameters are as $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(8,2,4,2,2)$. To reach the same spectral efficiency levels in the classical AFDM, the number of subcarriers is set to $8$, and QPSK and $8$-PSK are utilized correspondingly. The both schemes do not satisfy the full diversity order condition. Specifically, the proposed AFDM-PIM does not satisfy Condition 1 in Theorem 3, while the AFDM does not satisfy Theorem 1 of [17]. It can be seen from Fig. 5 that the AFDM-PIM demonstrates an about $2$ dB gain in the SNR compared to the AFDM scheme at the BER level of $10^{-3}$. This is because the AFDM-PIM allows lower order constellations to reach the same spectral efficiency as the AFDM. The results of Fig. 5 indicates that the AFDM-PIM offers a viable alternative for communication under doubly dispersive channels.
Figure 6: Performance comparisons between the proposed AFDM-PIM and the AFDM-IM scheme given two different spectral efficiencies.
In Fig. 6, we evaluate the BER performance of the proposed AFDM-PIM and the AFDM-IM [29] under a doubly dispersive channel with the number of paths $P\!=\!3$ and given two spectral efficiencies of $2$ and $3$ bit/s/Hz. For the proposed AFDM-PIM, $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})$ are set to $(8,2,4,1,2)$, and BPSK and QPSK are employed to reach the spectral efficiencies of $2$ and $3$ bit/s/Hz, respectively. For the AFDM-IM scheme, each group comprises $n$ subcarriers, with $a$ subcarriers activated at each transmission. To reach the same spectral efficiency levels for the AFDM-IM, $(n,a)\!=\!(4,2)$ and $(n,a)\!=\!(8,7)$ are employed, respectively, and $8$-PSK is utilized. Note that the both schemes do not meet the full diversity order condition. It can be seen from Fig. 6 that the AFDM-PIM scheme yields better BER performance than the AFDM-IM scheme, with an SNR gain of about $2$ dB at the BER level of $10^{-3}$. The enhanced performance is attributable to the capability of the AFDM-PIM to employ lower order constellations than the AFDM-PIM, while reaching the same spectral efficiency as the AFDM-IM.
Figure 7: Comparison of the theoretical ABEP upper bound and the simulated BER of the proposed AFDM-PIM given two different Doppler shifts.
VI-C Accuracy of PEP-based Theoretical Analysis
Fig. 7 presents a comparison of the theoretical ABEP upper bound and the simulated BER of the proposed AFDM-PIM scheme over a doubly dispersive channel with the number of paths $P\!\!=3$. The spectral efficiency of 1.5 bit/s/Hz is considered, $(N,G,\lambda,d_{\mathrm{max}})$ are set to $(4,2,2,0)$ and BPSK is employed as the modulation scheme. Besides, the maximum normalized Doppler shifts are set to $\alpha_{\mathrm{max}}\!=\!1$ and $\alpha_{\mathrm{max}}\!=\!2$, corresponding to the high-speed and ultra high-speed scenarios, respectively. It can be seen from Fig. 7 that in the both cases, the theoretical ABEP results deviate from the simulated results in the low-SNR region. This is because the theoretical ABEP upper bound is subject to several approximations, which inevitably becomes inaccurate when the noise is dominant. On the other hand, the theoretical results closely match the simulated BER curves at sufficiently high SNRs, which demonstrates the validity of the PEP analysis for the AFDM-PIM presented in Subsection IV-A.
VII Conclusions
In this paper, we have proposed the novel AFDM-PIM scheme to combat time-frequency doubly selective channel fading. In our proposed scheme, the pre-chirp parameters on subcarriers are no longer fixed but are selected from a predefined alphabet. This enables additional bits to be transmitted through indexing the specific pre-chirp parameter values on subcarriers, thereby enhancing both spectrum and energy efficiency. The PEP-based theoretical BER upper bound for the proposed AFDM-PIM scheme with the ML detection has been derived and verified through simulations, and the full diversity order conditions of the proposed AFDM-PIM under doubly dispersive channels have been derived. Furthermore, we have presented a PSO-based optimization algorithm to design the pre-chirp parameter alphabet. Both analytical and simulation results have demonstrated that the proposed AFDM-PIM exhibits enhanced spectral efficiency and superior error performance in comparison to classical multi-carrier modulation schemes.
-A Proof of Theorem 3
Proof.
First, we show that Condition 1 is necessary for the AFDM-PIM to achieve the full diversity order. Assume that Condition 1 is not met, i.e.,
$P>(d_{\mathrm{max}}+1)(2\alpha_{\mathrm{max}}+1).$ (62)
Under this assumption, according to the definition of $\mathrm{loc}_{p}\!=\!(\alpha_{p}+2Nc_{1}d_{p})_{N}$, the following situation must be true
$\exists a,b\in[1,\ldots,P],a\neq b,~{}\text{such that}~{}\mathrm{loc}_{a}=\mathrm{loc}_{b}.$ (63)
The corresponding two columns in the matrix $\mathbf{\Phi}(\delta)$ can be expressed as (64) and (65) at the top of the next page, where $\hat{H}_{(\cdot)}[\cdot]\hat{x}_{(\cdot)}$ and $H_{(\cdot)}[\cdot]x_{(\cdot)}$ represent the corresponding elements in $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ and $\mathbf{\Phi}(\mathbf{x})$, respectively.
$\begin{bmatrix}\hat{H}_{a}[0,\mathrm{loc}_{a}]\hat{x}_{\mathrm{loc}_{a}}-H_{a}[0,\mathrm{loc}_{a}]x_{\mathrm{loc}_{a}}\\ \hat{H}_{a}[1,(\mathrm{loc}_{a}+1)_{N}]\hat{x}_{(\mathrm{loc}_{a}+1)_{N}}-H_{a}[1,(\mathrm{loc}_{a}+1)_{N}]x_{(\mathrm{loc}_{a}+1)_{N}}\\ \vdots\\ \hat{H}_{a}[N-1,(\mathrm{loc}_{a}+N-1)_{N}]\hat{x}_{(\mathrm{loc}_{a}+N-1)_{N}}-H_{a}[N-1,(\mathrm{loc}_{a}+N-1)_{N}]x_{(\mathrm{loc}_{a}+N-1)_{N}}\end{bmatrix},$ (64)
$\begin{bmatrix}\hat{H}_{b}[0,\mathrm{loc}_{b}]\hat{x}_{\mathrm{loc}_{b}}-H_{b}[0,\mathrm{loc}_{b}]x_{\mathrm{loc}_{b}}\\ \hat{H}_{b}[1,(\mathrm{loc}_{b}+1)_{N}]\hat{x}_{(\mathrm{loc}_{b}+1)_{N}}-H_{b}[1,(\mathrm{loc}_{b}+1)_{N}]x_{(\mathrm{loc}_{b}+1)_{N}}\\ \vdots\\ \hat{H}_{b}[N-1,(\mathrm{loc}_{b}+N-1)_{N}]\hat{x}_{(\mathrm{loc}_{b}+N-1)_{N}}-H_{b}[N-1,(\mathrm{loc}_{b}+N-1)_{N}]x_{(\mathrm{loc}_{b}+N-1)_{N}}\end{bmatrix},$ (65)
$\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})-\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}=\sum_{p=1}^{P}\sum_{n=1}^{N}\left|H_{p}^{\prime}\left[n,\operatorname{loc}_{p}+n\right]x^{\prime}_{\operatorname{loc}_{p}+n}-H_{p}\left[n,\operatorname{loc}_{p}+n\right]x_{\operatorname{loc}_{p}+n}\right|^{2}$
$\hskip 28.45274pt=\sum_{p=1}^{P}\sum_{n=1}^{N}\left|e^{\textsf{j}\frac{2\pi}{N}\left[Nc_{2,\operatorname{loc}_{p}+n}^{\prime}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}^{\prime}n^{2}\right]}x^{\prime}_{\operatorname{loc}_{p}+n}]-e^{\textsf{j}\frac{2\pi}{{N}}\left[{N}c_{2,\operatorname{loc}_{p}+n}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}n^{2}\right]}x_{\operatorname{loc}_{p}+n}]\right|^{2}\!\!.\!$ (74)
Based on (III-D), the positions of the non-zero entries in the matrices $\mathbf{H}_{a},\hat{\mathbf{H}}_{a},\mathbf{H}_{b}$ and $\hat{\mathbf{H}}_{b}$ are consistent. In the instances, such as when both $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ and $\mathbf{\Phi}(\mathbf{x})$ contain only a single non-zero element at the same position, (64) and (65) are linearly correlated. Therefore, $\mathbf{\Phi}(\delta)$ cannot be full rank, i.e., the assumption (62) is false.
Besides, when $\mathbf{\Phi}(\delta)$ achieves the full diversity order, the channel paths with different delay values or distinct Doppler frequency shifts are distinguished within the DAFT domain, as illustrated in Fig. 3, i.e., $(d_{\mathrm{max}}+1)(2\alpha_{\mathrm{max}}+1)\leq N$. Therefore, Condition 1 is a prerequisite for achieving the full diversity order.
[…]
Given constants $\ell_{n,m}$, $\hat{H}_{p}\left[n,\operatorname{loc}_{p}+n\right]$ can be expressed as $\hat{H}_{p}[n,\operatorname{loc}_{p}+n]=\ell_{n,m}H_{p}[n,\operatorname{loc}_{p}+n]$. According to (III-D), the entries of $\bm{\gamma}_{p}$ are given by
$\bm{\gamma}_{p}[n]=$ $H_{p}[n,\mathrm{loc}_{p}+n]{\rho}_{n,m}$
$=$ $e^{\textsf{j}2\pi\left[c_{2,\operatorname{loc}_{p}+n}\left(\operatorname{loc}_{p}+n\right)^{2}-c_{2,n}n^{2}\right]}$
×$\times e^{\textsf{j}\frac{2\pi}{N}[-(\operatorname{loc}_{p}+n)+c_{1}d_{p}^{2}]}{\rho}_{n,m},$ (66)
where ${\rho}_{n,m}=\left(x_{\mathrm{loc}_{p}+n}-\ell_{m,n}\hat{x}_{\mathrm{loc}_{p}+n}\right)$. Now assume that a set of numbers $\{\beta_{p}\}_{p=1}^{P}$ that are not all zero satisfy
$\sum_{p=1}^{P}\beta_{p}\bm{\gamma}_{p}=\mathbf{0}.$ (67)
Without loss of generality, $\beta_{b}$ is assumed to be the non-zero number, $b\!\in\!\{1,2,\ldots,P\}$. The relationship for any index $n$ can be obtained as
$\sum_{p=1}^{P}\beta_{p}\bm{\gamma}_{p}[n]=\sum_{p=1}^{P}\frac{\beta_{p}}{\beta_{b}}\bm{\gamma}_{p}[n]=0,\quad\forall n.$ (68)
From (68), it is not difficult to obtain
${\rho}_{n,b}=-\!\!\!\sum_{p=1,p\neq b}^{P}\frac{\beta_{p}}{\beta_{b}}\frac{\bm{\gamma}_{p}[n]}{H_{p}[n,\mathrm{loc}_{p}+n]}$
$=-e^{-\textsf{j}2\pi c_{2,\mathrm{loc}_{b}}\mathrm{loc}_{b}^{2}}e^{\textsf{j}\frac{2\pi}{N}(Nc_{1}(-d_{b}^{2}+\mathrm{loc}_{b}d_{b}))}$
×$\hskip 5.69054pt\times\!\!\!\sum_{p=1,p\neq b}^{P}\!\!\!e^{\textsf{j}2\pi\big{(}c_{2,\mathrm{loc}_{p}+n}{(\mathrm{loc}_{p}+n)}^{2}+c_{1}d_{p}^{2}-\frac{\mathrm{loc}_{p}d_{p}}{N}\big{)}}\frac{\beta_{p}}{\beta_{b}}{\rho}_{n,m}.$ (69)
Since ${\rho}_{0,b}$ is a signal error term unrelated to $c_{2,n}$, the phase of $\frac{\beta_{p}}{\beta_{b}}(m\neq b)$ must contain $2\pi\big{(}c_{2,\mathrm{loc}_{b}}\mathrm{loc}_{b}^{2}-c_{2,\mathrm{loc}_{p}+n}(\mathrm{loc}_{p}+n)^{2}\big{)}$ to eliminate the influence of irrational numbers. On the other hand, if (-A) holds, there exists another non-zero $\beta_{a}$ ($a\neq b$). Similar to the derivation process of (-A), the phase of $\frac{\beta_{p}}{\beta_{a}}(m\neq a)$ must contain $2\pi\big{(}c_{2,\mathrm{loc}_{a}}\mathrm{loc}_{a}^{2}-c_{2,\mathrm{loc}_{p}+n}(\mathrm{loc}_{p}+n)^{2}\big{)}$. Hence, the relationship between $\frac{\beta_{a}}{\beta_{b}}$ and $\frac{\beta_{b}}{\beta_{a}}$ can be obtained as $\frac{\beta_{a}}{\beta_{b}}\cdot\frac{\beta_{b}}{\beta_{a}}=1$, i.e.,
$e^{\textsf{j}2\pi c_{2,\mathrm{loc}_{b}}\mathrm{loc}_{b}^{2}}e^{-\textsf{j}2\pi c_{2,\mathrm{loc}_{a}+n}(\mathrm{loc}_{a}+n)^{2}}$
×$\hskip 11.38109pt\times e^{\textsf{j}2\pi c_{2,\mathrm{loc}_{a}}\mathrm{loc}_{a}^{2}}e^{-\textsf{j}2\pi c_{2,\mathrm{loc}_{a}+n}(\mathrm{loc}_{a}+n)^{2}}\vartheta=1,~{}\forall n,$ (70)
where $\vartheta$ is a complex number whose phase does not contain $c_{2,n}$. Given that all the $c_{2,n}$ are irrational numbers, the phase on the left-hand side of (-A) cannot be an integer multiplying $2\pi$. This implies that the imaginary part is not zero, and (-A) does not hold. Consequently, the assumption of (67) is invalid, which means that the column vectors of the matrix $\mathbf{\Phi}(\delta)$ are linearly independent, i.e., the rank of $\mathbf{\Phi}(\delta)$ is $P$. Combining this result with the definition of diversity order (44), it is concluded that the AFDM-PIM can achieve the full diversity order. ∎
-B Derivation of $O_{k,j}(\mathcal{P}_{c})$ (49)
Substituting $\mathbf{\Phi}_{j}^{r}(\mathbf{x})\!=\![\mathbf{H}_{1}\mathbf{x},\ldots,\mathbf{H}_{P}\mathbf{x}]$ and $\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})\!=\![\mathbf{H}_{1}^{\prime}\mathbf{x}^{\prime},\ldots,\mathbf{H}_{P}^{\prime}\mathbf{x}^{\prime}]$ into (47) leads to
$O_{k,j}(\mathcal{P}_{c})=\sum_{\mathbf{x}^{\prime}\!,\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\left\|\mathbf{H}_{p}^{\prime}\mathbf{x}^{\prime}-\mathbf{H}_{p}\mathbf{x}\right\|_{\mathrm{2}}.$ (71)
According to (III-D), the elements of $\mathbf{H}_{p}\mathbf{x}$ can be expressed as
$\mathbf{H}_{p}\mathbf{x}[n]=$ $H_{p}\left[n,\operatorname{loc}_{p}+n\right]x_{\mathrm{loc}_{p}+n}$
$=$ $e^{\textsf{j}2\pi\big{(}c_{2,\operatorname{loc}_{p}+n}(\operatorname{loc}_{p}+n)^{2}-c_{2,n}n^{2}d_{p}\big{)}}$
×$\times e^{\textsf{j}\frac{2\pi}{N}\big{(}-(\operatorname{loc}_{p}+n)+c_{1}d_{p}^{2}\big{)}}x_{\mathrm{loc}_{p}+n}.$ (72)
Moreover, the elements of $\mathbf{H}_{p}^{\prime}\mathbf{x}^{\prime}-\mathbf{H}_{p}\mathbf{x}$ can be expressed as
$\left(\mathbf{H}_{p}^{\prime}\mathbf{x}^{\prime}-\mathbf{H}_{p}\mathbf{x}\right)[n]=H_{p}^{\prime}\left[n,\operatorname{loc}_{p}+n\right]x^{\prime}_{\operatorname{loc}_{p}+n}$
$\hskip 22.76219pt-H_{p}\left[n,\operatorname{loc}_{p}+n\right]x_{\operatorname{loc}_{p}+n}$
$\hskip 14.22636pt=\bigg{(}e^{\textsf{j}\frac{2\pi}{N}\big{(}Nc_{2,\operatorname{loc}_{p}+n}^{\prime}(\operatorname{loc}_{p}+n)^{2}-{N}c_{2,n}^{\prime}n^{2}\big{)}}x^{\prime}_{\operatorname{loc}_{p}+n}$
$\hskip 22.76219pt-e^{\textsf{j}\frac{2\pi}{{N}}\big{(}{N}c_{2,\operatorname{loc}_{p}+n}(\operatorname{loc}_{p}+n)^{2}-{N}c_{2,n}n^{2}\big{)}}x_{\operatorname{loc}_{p}+n}\bigg{)}$
×$\hskip 22.76219pt\times e^{\textsf{j}\frac{2\pi}{{N}}\big{(}-(\operatorname{loc}_{p}+n)d_{p}+{N}c_{1}d_{p}^{2}\big{)}}.$ (73)
Therefore, the norm of $\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})-\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}$ in (47) can be calculated as (-A), shown at the bottom of this page.
Utilizing the formula
$\mid a-b\mid^{2}=\mid a\mid^{2}-2\Re(ab^{*})+\mid b\mid^{2},$ (75)
where $a$ and $b$ are any complex numbers, (-A) can be expressed as
$\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})\!-\!\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}\!=\!\sum_{p=1}^{P}\sum_{n=1}^{N}\left|e^{\textsf{j}\theta_{n}^{\prime}}x^{\prime}_{\operatorname{loc}_{p}+n}\!-\!e^{\textsf{j}\theta_{n}}x_{\operatorname{loc}_{p}+n}\right|^{2}$
$\hskip 8.53581pt=\!\sum_{p=1}^{P}\sum_{n=1}^{N}\!2\!\left(1\!-\!\mathfrak{R}\left(x_{\operatorname{loc}_{p}+n}^{\prime}\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}e^{\textsf{j}(\theta_{n}^{\prime}-\theta_{n})}\right)\!\right)\!,\!$ (76)
where $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$ are the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, respectively, and $\theta_{n}^{\prime}$ and $\theta_{n}$ are given in (51).
According to Euler’s formula, we further simplify (-B) as
$\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})\!-\!\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}\!=\!\sum_{p=1}^{P}\!\sum_{n=1}^{N}\!1\!-\!\cos\left(\psi_{n}\!+\!\theta_{n}^{\prime}\!-\!\theta_{n}\right),$ (77)
where $\psi_{n}$ is defined in (50), i.e., it represents the phase difference between $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$. Substituting (77) into (47) leads to $O_{k,j}(\mathcal{P}_{c})$ of (49).